OpenStudy (anonymous):

Divide these polynomials. (x4 + x3 - 7x2 - 2x + 3) (x + 3) A. x3 - 3x2 + 2x + 2 B. x4 + 2x3 - x2 + x C. x3 + 2x2 - x + 1 D. x3 - 2x2 - x + 1

OpenStudy (anonymous):

Do long div like the last one: 3x^2 + 2x + 1 ________________________ 3x+1 ) 9x^3 + 9x^2 + 5x +1 - 9x^3 + 3x^2 -------------- 6x^2 + 5x - 6x^2 + 2x ----------- 3x + 1 - 3x + 1 --------- 0

OpenStudy (anonymous):

\[ \frac{x^4+x^3-7 x^2-2 x+3}{x+3}=\frac { (x+3)(x^3-2 x^2-x+1) } {x+3)}=x^3-2 x^2-x+1 \]